OpenStudy (nali):

PLZ HELP 3x+7y=12 6x-ky=6 for which of the following values for k will the system of equations above has no solution ? -14 -7 0 7 14

OpenStudy (nali):

@jim_thompson5910 plz help

OpenStudy (hawkfalcon):

Did you try plugin them into a calculator?

OpenStudy (nali):

i dont get ur answer @hawkfalcon

OpenStudy (nali):

what do u mean by replace?

OpenStudy (lopus):

you can replace x in 9x+(7-k)y=18

OpenStudy (nali):

how did u know the value of x?

OpenStudy (lopus):

you plus two equations andsolve for x

OpenStudy (nali):

sorry i do not get u can u plz show ur work

OpenStudy (nali):

@lopus can u plz show step by step i get u so far but how ,in the end,did u find the value of ? i need to learn how to do the problem

OpenStudy (nali):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Notice how 3x/6x = 1/2 So for the system to be inconsistent, then 7y/(-ky) = 1/2 should be true as well

OpenStudy (nali):

ok so is it fourteen ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Negative 14. You lost a sign somewhere.

OpenStudy (nali):

oh yeah thanku so much :)

jimthompson5910 (jim_thompson5910):

sure thign

jimthompson5910 (jim_thompson5910):

thing*

jimthompson5910 (jim_thompson5910):

no, it's -14

jimthompson5910 (jim_thompson5910):

6x-ky=6 6x-(-14)y=6 6x+14y = 6

jimthompson5910 (jim_thompson5910):

no it won't 3x+7y=12 6x+14y = 6 has no solutions

OpenStudy (lopus):

ok im write wrong 3x-7y sorry

jimthompson5910 (jim_thompson5910):

np